Assume an isotropic antenna operating at 10 GHz, find the effective area and antenna gain.
1
Expert's answer
2021-09-21T02:06:09-0400

The effective area 

"A_e=k \frac{\pi D^2}{4}\\
46=10 \log_{10} k(\frac{\pi D^2}{\lambda})^2\\
\frac{\pi D^2}{\lambda}=269.04\\
D=\frac{269.04* \lambda}{\pi}=\frac{269.04* 0.025}{3.14}=2.14\\
\therefore A_e=k \frac{\pi D^2}{4}= \frac{0.55*3.14*2.14^2}{4}=1.98 m^2\\"


Antenna gain

"G=10 \log_{10} k(\frac{\pi D^2}{\lambda})^2\\
G=10 \log_{10} *0.55*(\frac{3.14* 30}{0.075})^2\\
G=59.38dB"